OpenStudy (anonymous):

What is the length of the focal width? x^2=12y

OpenStudy (anonymous):

length of focal width for the parabola x^2=4by is 4b so for x^2=4*3*y it should be...

OpenStudy (anonymous):

are you asking me to finish that.. lol

OpenStudy (anonymous):

well.. idk